OpenStudy (anonymous):

y<4x-7 y>x-4 (the arrow facing > has a line under it) A(4, 0) B(6, -2) C(8, 6) D(0, -4)

OpenStudy (texaschic101):

lets sub and find our answer...keeping in mind, what we sub in has to satisfy both inequalities or it won't be a solution. y < 4x - 7 and y >= x - 4 (4,0)...x = 4 and y = 0 now we sub 0 < 4(4) - 7 0 < 16 - 7 0 < 9 (correct) now check the other inequality y >= x - 4 0 >= 4 - 4 0 >= 0 (correct) wow...that just happened to be the first one. (4,0) is your solution to these inequalities